March 18-21

– Three stores provided ~44 hours of luminosity
– Time Line Generator problems fixed
– Tevatron RF station trip caused luminosity loss
– Controls personnel replaced bad Meson Septa motors
– Booster personnel repaired chopper
– Lightning strikes caused store 8594 to abort and many other accelerator problems
– Tevatron cryo magnet failed and must be replaced – Tevatron off for 10-11 days

* The Integrated Luminosity for the period from 3/14/11 to 3/21/11 was 70.75 inverse picobarns. NuMI is off for target replacement.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts